# Learning From Software Failures: A Case Study at a National Space Research Center

### v2

@@ 5.1.1. Gathering lessons from failures is informal
Across all subjects, there was no structured or consistent process for gathering lessons learned from failures; instead, such efforts were ad hoc, informal, and dependent on personal initiative. All subjects stated that formal processes for reflection were absent. As S5 noted, “I don’t think that we have a formal process on how to make sure that [failures] doesn’t happen again.” Others echoed this absence, with S6 stating, “We don’t have a very strict process…not really,” and S3 explaining that group reflection is not part of their routine: “There’s not a big reflective period where we say, ‘Okay, this error happened…how should we do it in the future?”’ Even recalling a single such meeting proved difficult for some; S5 admitted, “I can’t think of a single time where we just sat there and discuss…issue to avoid it in the future.” S1 emphasized “we study the problem in order to fix the problem and not…to retrieve some knowledge out of it.” Instead, participants described informal and improvised efforts to reflect on failures. These were typically team-driven and irregular, such as a one-time feedback meeting described by S8: “We’re going to do this actually in a couple of weeks…but it’s not standard.” S9 confirms, “We do it…on the fly…But negative point is of course that there’s no real structure, it’s more like sharing experiences.”
@@ 5.1.2. Documentation of lessons learned is inconsistent & fragmented
In the absence of formal processes, subjects rely on tacit knowledge rather than accessible records. When asked about lessons learned, S10 stated, “Not really lessons learned, no. I think they’re more like internally…we learned from it.” Likewise, S8 emphasized the implicit nature of such knowledge: “Not explicit, but implicit…you learn how to do it in a better way, but this is just internal knowledge.” S2, S6, and S10 shared examples where even when recurring issues are discussed, the knowledge is rarely written down, “I don’t think we document it anywhere to be honest…Maybe it would be a good idea.”
@@ 5.1.3. Lessons learned are shared & applied informally
  • A couple situations…where issues…were recurring…[and often] someone else on the team…[would recollect], ‘Oh yeah, I remember I faced this too. You have to do…’ So that’s probably the worst case…to handle those things, but that’s how it’s done…You have to hope…[to find] someone [who remembers that they] already experienced it…there’s not a process to look it up.
@@
To enable feedback cycles, S7 mentions questionnaires, though no other subject was aware of them, and they remain unused. To enable feedback cycles, S7 mentioned questionnaires: one to reflect on failures, “a questionnaire to discuss…what’s wrong?…can this happen again?”, and another to apply lessons during, “design and getting…requirements…to get the criticality of system.” However, when we pointed out that no other subject referenced these, S7 admitted, “Normally you should use it always…But people don’t…we cannot [enforce] everyone to do it,” and as a result, “failures are made again and again.” As S2 said, “You’ll [have to] be self-motivated to search the wiki…GitLab, ask the seniors…otherwise…there’s no process to bring lessons down to you.”
@@ 5.1.4. Transferability
The pilot interviews matched the case study data: gathering, documenting, and applying lessons from failure were also ad hoc for those subjects.
Pilot participants confirmed that lessons from failures are gathered informally and often depend on personal motivation. PTeam described a one-off, “brainstorming workshop…where we discussed… production issues we have being observing…and not having those repeat again,” but this was an exception, not the norm. PS3 admitted, “I think I was the only one who would [reflect] such a thing.” Subjects explicitly stated the absence of formal reflection processes, with S5’s acknowledgment that PS1 “I don’t think we have a formal process to make sure that [failures] doesn’t happen again.” PS1 echoed, “There’s no process for that.”
Documentation practices in the pilot study mirrored the case study’s inconsistencies, due to time and resource constraints. PS1 explained “we’re very short-staffed…we don’t have time to write new documentation.” PS2 noted that postmortems are recorded simply as meeting minutes without structure: “there isn’t a specific template for the postmortem data.” PS1 said information was recorded in a wiki, and echoed difficulty in finding useful content: “you can search around…but it’s hard to actually get stuff you can sit down and read.”
Application of lessons from failures is limited, and their transfer also relied heavily on informal mentorship. PS3 and PS1 explicitly stated lessons were “not referenced once [the project] is closed.” Knowledge transfer was largely passive and junior engineers were expected to take initiative: “They don’t really incentivize that…it’s more on the younger engineers to ask for it” (PS1). Formal training on past failures was absent, with one participant noting, “it’s out there, but it’s not specifically addressed” (PS4), and new hires typically learned by observation: “they just watch how other team members are responding” (PTeam).
@@ 5.2.2. Confidence in general practices undermines the perceived need for learning from failures
However, some participants acknowledged limitations in relying solely on best practices. S7 criticized the superficiality of standards, “with the standards…the common agreement is…so weak…It didn’t help you normally…[because] they don’t share too much knowledge,” consistent with prior findings (Chen et al., 2024). S1 noted another drawback, “[It’s] a 150–200 page PDF with a lot of rules…even I can’t tell you all the rules. [We] should know where this is, but not learned by heart…[it’s] not [referenced] that often.” Consequently, the HRO standard they sometimes follow requires lessons learned to be gathered, but they do not seem to do it consistently, as S2 shares, “The standard from the…Space Agency…[requires us to] collect lesson learned…but this doesn’t apply for all of our projects.”
@@
Notably, some subjects attributed recurring failures to the inexperience of junior engineers — masking systemic gaps in how failure knowledge is transferred. S7 observed, “I see [failures] happened again and again because…young software engineers…didn’t have the experience and the sharing of experiences is limited.” They defended their on-boarding process, claiming it mitigates these gaps: “There aren’t really gaps, there’s training. The only issue we have is time to get the lessons and [applying] the lessons. That’s why we have an on‑boarding process.” Responsibility for avoiding recurring failures was shifted to individuals, particularly newcomers, rather than organizational processes to learn from failures. S7 explained, “as long as the process is understandable for new employees…for experienced engineers it might not be necessary…[they already have] a big knowledge base.” This assumption risks perpetuating knowledge silos and obscuring structural challenges in failure knowledge transfer.
5.2.3. Ad hoc learning sometimes prevents recurrence—but is rare
@@
In a few cases, serious or recurring failures prompted technical or process improvements, though these were typically driven by impactful incidents and individual initiative rather than structured processes. S7 described a recurring, “big fail[ure]…the project manager called and I hear…crying and…[he says,] ‘everything is not working…we cannot start the satellite…nothing works’…[I respond], ‘Yeah, I know this failure, this was fixed three months ago’…[Our] lessons learned…[from this failure led to] this modeling…group…and test process…[and] a research project.” In another case, S6 explained how a post-mission reflection led to improvements: “From results of mission one…we learn a lot… [and] we are applying it for the new mission.” In a different case, S2 recalled that after a recurring issue, the team collectively learned, documented a guideline, and shared it: “we…define a guideline…how to…and what not to do…to not get this problem,” and as a result the problem did not recur “because the people who joined afterwards, they follow the same instructions.” S3 noted that from a prior failure, they, “try to have a lot of checks beforehand…so later you don’t have as many pitfalls.”
@@ 5.2.4. Transferability
Pilot subjects agreed that recurring issues remain a problem due to the lack of structured processes for failure learning. PS4 admitted, “we come across the same issues over and over…it gets worse and worse”, echoing the case study’s examples of unresolved issues resurfacing across projects. Pilot subjects described formal structures like root cause analysis and documented action items. However, PS4 noted, “I don’t know if anybody really makes sure the action items get done”, and PS3 added, “they just talked about it and left the meeting”. This lack of a structured feedback-loop perpetuates the illusion of effective failure management.
@@ 5.3. Theme 3: Challenges within the learning-from-failures process
Themes 1 and 2 highlighted the ad hoc nature of learning from failures. Since many subjects were aware of the ill effects, we examine why more robust processes have not taken root.
@@ 5.3.2. Informal & flexible SDLC limits learning structures
Many acknowledged that other engineering disciplines within the center, particularly hardware teams, follow more rigorous post-failure processes due to the high cost of mistakes. As S5 observed, “If you do something stupid…in software then you just update your software…but in hardware, you burn money…because it’s harder to fix in the end.” S4 added: “So…they have to move more careful than software people do”, as did S1, “And I think for this reason there is probably a process.” As a result, S1 notes, “In other institutes…doing satellite hardware, they have knowledge base on what are typical faults or failures…in our institute there’s nothing like this.” This believe that software failures are less expensive may reduce pressure to institutionalize learning. S6 contradicts the other subject by noting the high cost of software fixes after deployment (Zage and Zage, 2003; McConnell, 2001; Stecklein et al., 2004), “even with…a software update…[on] something flying on the [space station], you need to…[do] software approval and this can take…half a year to a year…[Even if] a software update..[is] ‘free’…it’s not easy and you want to get the software right the first time.”
@@ 5.3.3. Knowledge loss from team turnover
S5 reinforced this need, “considering…that people are leaving quite frequently and their knowledge is lost…something like lessons learned would definitely be helpful.” When such documentation does exits, S4 recalls its usefulness, “I have used what other team members have written [about issues]…it saved me a lot of time…especially in cases [where] more people are joining the team or you could also leave at some point.”
@@ 5.3.5. Transferability
The pilot study reinforces the case study’s findings that time and resource constraints severely limit structured failure learning. Pilot subjects cited lack of time as a major barrier. PS1 stated “[we] don’t have time to sit down and write everything that [we’ve] learned because [we’re] so busy.” PTeam added, “with agile, it is hard to keep up with both documentation and development. Time and priorities are constraints.”
@@ 6. Discussion and Future Work
This section reflects on whether current practices enable effective learning from failures, considers what broader software engineering practices could offer this HRO, highlights strengths of this HRO’s practices, and outlines future directions.
@@ 6.1. Learning from failure: Are we equipped?
This study set out to understand how software engineers learn from failures, with the assumption that HROs, such as a national space research center, would exemplify best practices (cf. section 4.2). We selected this case study precisely because we expected its teams to operationalize the structured, cyclic processes for capturing, disseminating, and applying failure knowledge that are prescribed in standards for high-reliability engineering (Christianson et al., 2011). However, our findings challenge this assumption. Prior work proposed structured techniques for gathering, documenting, and sharing lessons from failures, with limited guidance on applying them (cf. section 2.2). In contrast, practices at this HRO were largely ad hoc: recommended techniques for gathering, documenting, and sharing were rarely followed, and offered little to learn about systematic application. While prior literature has recommended various structured practices for gathering, documenting, sharing, and applying lessons from failures (cf. section 2.2), practices at this HRO were ad-hoc. Despite operating in a high-stakes, reliability-critical environment, engineers at this HRO described largely informal, ad hoc, and inconsistent approaches to learning from software failures (cf. section 5.1). Many participants acknowledged recurring failures that were not systematically addressed, with lessons learned often confined to individual memory, undocumented conversations, or informal mentorship (cf. section 4.2). This pattern extended across our pilot study as well, which included subjects from other HROs and sectors (cf. section 5.1.4section 5.2.4).
@@
The challenges faced by this organization to learn from failures mirror barriers in conducting postmortems identified two decades ago (Kasi et al., 2008) (cf. section 5). The persistence of these long-standing obstacles, even in a mission-critical HRO, underscores how deeply entrenched such barriers are in software engineering practice. These findings raise a foundational question: If even such mission-critical HROs lack structured learning-from-failure practices, where should we look for models of effective failure learning? And a broader concern: even though failure analysis is widely recommended in both safety-critical systems and modern DevOps cultures, do software teams actually know when, why, and how to learn from failures effectively? Our results suggest that even in organizations theoretically aligned with the principles of continuous improvement and resilience, failure learning remains underdeveloped, reactive, and fragile. This disconnect points to a pressing need: the software engineering field may require clearer structures, practices, and cultural reinforcements to support failure learning. Without these, even well-intentioned teams may repeat the same mistakes, not due to negligence, but because they lack processes or tools for turning failures into actionable knowledge.
@@
Future work could look to domains where incentives for avoiding recurring failures are more tightly coupled to financial, contractual, or regulatory obligations. For example, cloud infrastructure providers with stringent service level agreements (Beyer et al., 2016), banking systems with financial repercussions and regulatory oversight (Pri, 2021), or government agencies with centralized engineering governance (NAS, [n. d.]a; Les, 2024). These settings may offer more developed models for learning from failures, due to their financial, contractual, or regulatory pressures.
@@ 6.2. What could broader software engineering practices teach this HRO?
While our case site operates in a mission-critical domain, many of its failure reflection and documentation practices lag behind those seen in large IT organizations. Postmortem practices at some of these organizations, such as structured templates and searchable incident databases, are well documented in practice and literature (Beyer et al., 2016; Allspaw, 2012). In contrast, subjects in our study described ad hoc gathering of lessons and fragmented documentation (cf. section 5.1).
Our results suggest that HROs could benefit from adapting practices from IT organizations, particularly in how they document and track incidents. Structured postmortem templates, searchable failure databases, and lightweight regular reflection could offer scalable processes to support failure learning without introducing excessive bureaucracy. Formalizing mentorship systems, on-boarding materials, or guidelines for recurring failure modes, could further strengthen resilience against knowledge loss.
Notably — as both the case study and pilot interviews revealed — even when lessons are documented, their application and integration into the SDLC was inconsistent. Similarly, even in large IT and DevOps organizations, where postmortems and tooling are sometimes adopted, it remains unclear whether these efforts lead to meaningful learning or sustained changes in their SDLC. This raises an important question: How do, and how can, learning-from-failure translate into sustained improvements in the SDLC? The feedback cycle from lessons to tangible process change remains poorly understood and, consequently, poorly supported by researchers.

### v3

@@ 5.1.1. Gathering lessons from failures is informal
Across all subjects, there was no structured or consistent process for gathering lessons learned from failures; instead, such efforts were ad hoc, informal, and dependent on personal initiative. All subjects stated that formal processes for reflection were absent. As S5 noted, “I don’t think that we have a formal process on how to make sure that [failures] doesn’t happen again.” Others echoed this absence, with S6 stating, “We don’t have a very strict process…not really,” and S3 explaining that group reflection is not part of their routine: “There’s not a big reflective period where we say, ‘Okay, this error happened…how should we do it in the future?”’ S1 emphasized “we study the problem in order to fix the problem and not…to retrieve some knowledge out of it.” Instead, participants described informal and improvised efforts to reflect on failures. These were typically team-driven and irregular, such as a one-time feedback meeting described by S8: “We’re going to do this actually in a couple of weeks…but it’s not standard.” S9 confirms, “We do it…on the fly…But negative point is of course that there’s no real structure, it’s more like sharing experiences.”
@@ 5.1.2. Documentation of lessons learned is inconsistent & fragmented
In absence of formal processes, subjects rely on tacit knowledge rather than accessible records. When asked about lessons learned, S10 stated, “Not really lessons learned, no. I think they’re more internal…we learned from it.” Likewise, S8 emphasized the implicit nature of such knowledge: “Not explicit, but implicit…you learn how to do it a better way, but this is just internal knowledge.” S2, S6, and S10 shared examples where even when recurring issues are discussed, the knowledge is rarely written down, “I don’t think we document it anywhere to be honest…Maybe it would be a good idea.”
@@ 5.1.3. Lessons learned are shared & applied informally
  • A couple [of] situations…where issues…were recurring…[and often] someone else on the team…[would recollect], ‘Oh yeah, I remember I faced this too. You have to do…’ So that’s probably the worst case…to handle those things, but that’s how it’s done…You have to hope…[to find] someone [who remembers that they] already experienced it…there’s not a process to look it up.
@@
To enable feedback cycles, S7 mentioned questionnaires: one to reflect on failures, “a questionnaire to discuss…what’s wrong?…can this happen again?”, and another to apply lessons during, “design and getting…requirements…to get the criticality of system.” However, when we pointed out that no other subject referenced these, S7 admitted, “Normally you should use it always…But people don’t…we cannot [enforce] everyone to do it,” and as a result, “failures are made again and again.” As S2 said, “You’ll [have to] be self-motivated to search the wiki…GitLab, ask the seniors…otherwise…there’s no process to bring lessons down to you.”
@@ 5.1.4. Transferability
The pilot interviews matched the case study data: gathering, documenting, sharing, and applying lessons from failures were ad-hoc.
Pilot subjects shared that failure knowledge was gathered informally and often depended on personal motivation (cf. section 5.1.1). For example, the PTeam described a one-off, “brainstorming workshop…where we discussed…production issues we have being observing…and not having those repeat again,” but noted this was an exception, not the norm. PS3 admitted, “I think I was the only one who would [reflect].” Subjects stated the absence of formal reflection practices, with PS1 acknowledging, “we don’t have any kind of formal investigation process. It’s more, ‘just go fix it.”’
Documentation practices of failure knowledge in the pilot study mirrored the case study’s absence of structured practices, due to time and resource constraints (cf. section 5.1.2). PS1 explained “we’re very short-staffed…we don’t have time to write documentation.” PS2 noted that postmortems are simply recorded as “meeting minutes” without structure, since “there isn’t a specific template for the postmortem data,” and as a result, “I don’t believe we are at a point where we are collecting all that [postmortem] data.”
Application of lessons from failures was limited, and knowledge transfer often relied on informal mentorship (cf. section 5.1.3). PS1 reflected, “I’ve never heard anyone say, ‘how can we make sure [failures] don’t happen again?”’ PS2 stated that failure knowledge was “not referenced once [the project] is closed.” PS4 noted that, “action items [after failures] are put in a [wiki] or an email, so someone would manually track it…but I don’t know if anybody follows up and ensures that the action items get done.” PS1 explained that knowledge transfer was passive and relied on junior engineers taking initiative: “They don’t really incentivize…staff engineers to mentor [juniors]…to give them knowledge…it’s more on the younger engineers to ask for it.” S4 similarly remarked that lessons are “out there, but it’s not specifically addressed in training,” with the PTeam adding “they just watch how other team members are responding.”
@@ 5.2.2. Confidence in general practices undermines the perceived need for learning from failures
However, some participants acknowledged limitations in relying solely on best practices. S1 noted the overhead of following best practices from the organization’s software guideline, “[It’s] a 150–200 page PDF with a lot of rules…even I can’t tell you all the rules. [We] should know where this is, but not learned by heart…[it’s] not [referenced] that often.” Consequently, although the guideline requires them to gather lessons learned, they do not seem to do it consistently, as S2 shares, “The standard from the…Space Agency…[requires us to] collect lesson learned…but this doesn’t apply for all of our projects.” S7 criticized the superficiality of such guidance in describing how such processes should be implemented in practice: “With the standards…the common agreement is…so weak…It didn’t help you normally…[because] they don’t share too much knowledge,” consistent with prior findings (Chen et al., 2024).
@@
Notably, some subjects attributed recurring failures to the inexperience of junior engineers – masking systemic gaps in how failure knowledge is transferred. S7 observed, “I see [failures] happened again and again because…young software engineers…didn’t have the experience and the sharing of experiences is limited.” They defended their on-boarding process, claiming it mitigates these gaps: “There aren’t really gaps, there’s training. The only issue we have is time to get the lessons and [applying] the lessons. That’s why we have an on‑boarding process.” Responsibility for avoiding recurring failures was shifted to individuals, particularly newcomers, rather than organizational processes to learn from failures. S7 explained, “as long as the process is understandable for new employees…for experienced engineers it might not be necessary…[they already have] a big knowledge base.” This assumption risks perpetuating knowledge silos and obscuring structural challenges in failure knowledge transfer.
5.2.3. Ad hoc learning sometimes prevents recurrence–but is rare
In a few cases, serious or recurring failures prompted technical or process improvements, though these were typically driven by impactful incidents and individual initiative rather than structured processes. S7 described a recurring,
  • big fail[ure]. The project manager called and I hear…crying. [He says:] ‘we cannot start the satellite…nothing works.’
    [I reply:] ‘Yeah, I know this failure, it was fixed three months ago’.
    [From this failure, our] lessons learned was…this modeling group…a test process…[and] a research project.
@@ 5.2.3. Ad hoc learning sometimes prevents recurrence–but is rare
In another case, S6 explained how a post-mission reflection led to improvements: “From results of mission one…we learn a lot… [and] we are applying it for the new mission.” In a different case, S2 recalled that after a recurring issue, the team learned, documented a guideline, and shared it: “we…define a guideline…how to…and what not to do…to not get this problem,” and as a result the problem did not recur “because the people who joined afterwards followed the same instructions.” S3 noted that from a prior failure, they, “have a lot of checks beforehand…so later you don’t have as many pitfalls.”
@@ 5.2.4. Transferability
Pilot subjects also reported facing recurring issues due to lack of structured practices to learn from failures. PS4 admitted, “we have multiple root cause analyses with same [root cause], but we never prioritize getting it fixed…[so] we come across the same issues over and over…it gets worse and worse.” PS3 acknowledges the lack of emphasis to learn from failures, “they don’t care about failures. They just work past it and continue making failures.”
@@ 5.3. Theme 3: Challenges within the Learning-from-failures Process
Themes 1 and 2 highlighted the ad–hoc nature of learning from failures. Since many subjects were aware of the ill effects, we examine why more robust processes have not taken root.
@@ 5.3.2. Informal & flexible SDLC limits learning structures
Many acknowledged that other engineering disciplines within the center, particularly hardware teams, follow more rigorous post-failure processes due to the high cost of mistakes. As S5 observed, “If you do something stupid…in software then you just update your software…but in hardware, you burn money…because it’s harder to fix in the end.” S4 added: “So…they have to move more careful than software people do”, as did S1, “And I think for this reason there is probably a process.” As a result, S1 notes, “In other institutes…doing satellite hardware, they have [a] knowledge base on what are typical faults or failures…in our institute there’s nothing like this.” This believe that software failures are less expensive may reduce pressure to institutionalize learning. S6 contradicts the other subject by noting the high cost of software fixes after deployment (Zage and Zage, 2003; McConnell, 2001; Stecklein et al., 2004), “even with…a software update…[on] something flying on the [space station], you need to…[do] software approval and this can take…half a year to a year…[Even if] a software update..[is] ‘free’…it’s not easy and you want to get the software right the first time.”
@@ 5.3.3. Knowledge loss from team turnover
S5 reinforced this need, “considering…that people are leaving quite frequently and their knowledge is lost…something like lessons learned would definitely be helpful.” When such documentation does exits, S4 recalls its usefulness, “I have used what other team members have written [about issues]…it saved me a lot of time…especially in cases [where] more people are joining the team or…leaving.”
@@ 5.3.5. Transferability
Pilot study subjects shared challenges in learning from failures that mirrored those identified in the case study. All pilot subjects cited lack of time and resources as the major barrier (cf. section 5.3.1). PS1 explained, “trying to learn from failures…is very difficult to do…when you’re short staffed…and [we] don’t have time to sit down and write everything we’ve learned, because [we’re] so busy.” The PTeam similarly emphasized that “time and priorities are constraints.” PS1 expressed how this leads to knowledge loss during team turnover, “it becomes a big problem where senior engineers’ knowledge isn’t captured anywhere because they’re so short on time.” PS1 notes that even when failure knowledge is documented, it is often inaccessible,“you can search around…but it’s hard to actually get stuff you can sit down and read and get [lessons] out of.”
@@ 6.1. Learning from Failure: Are We Equipped?
This study set out to understand how software engineers learn from failures, with the assumption that HROs, such as a national space research center, would exemplify best practices (cf. section 4.2). We selected this case study precisely because we expected its teams to operationalize the structured, cyclic processes for capturing, disseminating, and applying failure knowledge that are prescribed in standards for high-reliability engineering (Christianson et al., 2011). However, our findings challenge this assumption. Prior work proposed structured techniques for gathering, documenting, and sharing lessons from failures, with limited guidance on applying them (cf. section 2.2). In contrast, despite operating in an HRO environment, our subjects described largely informal, ad-hoc, and inconsistent approaches to learning from software failures (cf. section 5.1). Many participants acknowledged recurring failures that were not systematically addressed, with lessons learned often confined to individual memory, undocumented conversations, or informal mentorship (cf. section 4.2). This pattern extended across our pilot study as well, which included subjects from other HROs and sectors (cf. section 5.1.4section 5.2.4).
@@
The challenges faced by this organization to learn from failures mirror barriers in conducting postmortems identified two decades ago (Kasi et al., 2008). The persistence of these long-standing obstacles, even in a mission-critical HRO, underscores how deeply entrenched such barriers are in software engineering practice. These findings raise a foundational question: If even such mission-critical HROs lack structured learning-from-failure practices, where should we look for models of effective failure learning? And a broader concern: even though failure analysis is widely recommended in both safety-critical systems and modern DevOps cultures, do software teams actually know when, why, and how to learn from failures effectively? Our results suggest that even in organizations theoretically aligned with the principles of continuous improvement and resilience, failure learning remains underdeveloped, reactive, and fragile. This disconnect points to a pressing need: the software engineering field may require clearer structures, practices, and cultural reinforcements to support failure learning. Without these, even well-intentioned teams may repeat the same mistakes, not due to negligence, but because they lack processes or tools for turning failures into actionable knowledge.
@@
Future work could look to domains where incentives for avoiding recurring failures are more tightly coupled to financial, contractual, or regulatory obligations. For example, cloud infrastructure providers with stringent service level agreements (Beyer et al., 2016), banking systems with financial repercussions and regulatory oversight (Pri, 2021), or government agencies with centralized engineering governance (NAS, [n. d.]a; Les, 2024), may offer more developed models for learning from failures.
@@ 6.2. What Could Broader Software Engineering Practices Teach this HRO?
While our case site operates in a mission-critical domain, many of its failure reflection and documentation practices lag behind those seen in large IT organizations. Postmortem practices at some of these organizations, such as structured templates (Dingsøyr, 2005; Beyer et al., 2016) and searchable incident databases (Beyer et al., 2016; NAS, [n. d.]b; Johnson, 2000), are well documented in practice and literature (Beyer et al., 2016; Allspaw, 2012). In contrast, subjects in our study described ad hoc gathering of lessons and fragmented documentation (cf. section 5.1).
Our results suggest that HROs could benefit from adapting practices from IT organizations, particularly in how they document and track incidents. Structured postmortem templates, searchable failure databases, and lightweight regular reflection could offer scalable processes to support failure learning without introducing excessive bureaucracy (Dingsøyr, 2005; Beyer et al., 2016), while also laying the foundational structure needed for future AI-supported tools to access and leverage failure knowledge at scale (cf. section 6.4). Formalizing mentorship systems, on-boarding materials, or guidelines for recurring failure modes, could further strengthen resilience against knowledge loss (Beyer et al., 2016).
Notably, as both the case study and pilot interviews revealed, even when lessons are documented, their application and integration into the SDLC was inconsistent. Similarly, even in large IT and DevOps organizations, where postmortems and tooling are sometimes adopted, it remains unclear whether these efforts lead to meaningful learning or sustained changes in their SDLC. This raises important questions: How do existing practices for learning from failures translate into improvements? And how can such practices be designed to ensure that lessons from failures lead to sustainable improvments? The feedback cycle from lessons to tangible improvements remains poorly understood and, consequently, poorly supported by researchers.